# Asymmetric Graph Error Control with Low Complexity in Causal Bandits

### v1

@@ Abstract
In this paper, the causal bandit problem is investigated, in which the objective is to select an optimal sequence of interventions on nodes in a causal graph. It is assumed that the graph is governed by linear structural equations; it is further assumed that both the causal topology and the distribution of interventions are unknown. By exploiting the causal relationships between the nodes whose signals contribute to the reward, interventions are optimized. First, based on the difference between the two types of graph identification errors (false positives and negatives), a causal graph learning method is proposed, which strongly reduces sample complexity relative to the prior art by learning sub-graphs. Under the assumption of Gaussian exogenous inputs and minimum-mean squared error weight estimation, a new uncertainty bound tailored to the causal bandit problem is derived. This uncertainty bound drives an upper confidence bound based intervention selection to optimize the reward. To cope with non-stationary bandits, a sub-graph change detection mechanism is proposed, with high sample efficiency. Numerical results compare the new methodology to existing schemes and show a substantial performance improvement in both stationary and non-stationary settings. Compared to existing approaches, the proposed scheme takes $67\%$ fewer samples to learn the causal structure and achieves an average reward gain of $85\%$.
@@
Causal bandit, linear structure equation model, graph identification, mutual information, upper confidence bound
@@ I Introduction
Investigation of cause-effect relationships is central to answering key questions in medicine, epidemiology, economics as well as the social and behavioral sciences (see e.g. [2, 3, 4]). Cause-effect relationships can be represented by Bayesian networks, in the form of directed acyclic graphs (DAGs). Identifying the causal structural model is essential to decision making and answering counterfactual questions.
@@
Current causal graph identification methods mainly fall into two categories: independence-based and score-based. Independence-based approaches, such as[2, 5], search for conditional independence over all possible graphs and thus computation complexity scales exponentially. To lower the complexity, recent score-based methods (see e.g. [6, 7, 8, 9]) formulate the structure learning problem as a continuous optimization task. In [7], an algorithm called GOLEM is proposed, where the likelihood score is combined with a soft DAG constraint to evaluate possible causal structures. The DAGMA scheme proposed in [8] characterizes acyclicity by a log-determinant function, with better-behaved gradients.
Decision-making or action-taking is natural within the context of causal inference. Despite this natural relationship, incorporation of causal methods within the context of multi-armed bandits (MAB) [10, 11], has only occurred recently. We note that separating graph identification from optimized interventions for MABs is inefficient as these two goals are coupled in causal MABs. Another feature not typically considered in the causal graph identification or causal MAB literature is the distinguishing between false positive and false negative errors in graph identification – we shall actively consider this feature.
@@
MABs provide a useful model for sequential decision-making problems, and as such have been employed in the design of clinical trials [12], recommendation systems [13], financial portfolio design [14], etc. In a MAB, an agent selects an action (arm) in each round and observes corresponding outcomes – the goal is to maximize the long-term cumulative reward. In the standard setting, the stochastic rewards generated by different arms are assumed to be statistically independent. To model realistic scenarios with dependence, causal bandits were introduced [11], where the causal structure is exploited to improve decision-making. DAGs can encode the causal relationship among factors that contribute to the reward [11]. Herein, we interpret the arms as different interventions on the nodes of a DAG and the reward as the stochastic outcome of a certain node. In the current work, we will provide new methods for both causal inference (graph identification) as well as the design of optimal interventions for causal MABs.
@@
Existing literature on causal bandits can be categorized based on their assumptions about the topology of the underlying DAG and the distribution of the interventions. While many works assume prior topology or DAG knowledge [11, 15, 16, 17], this is often not known in practice. In this work, knowledge of the underlying causal structure is not assumed.
@@
The setting without the knowledge of both topology and interventional distribution has been investigated recently [18, 19, 20]. An algorithm based on causal tree recovery is proposed in [18], with regret scaling logarithmically with the number of nodes. In [19], an auxiliary separating set algorithm is utilized to learn the causal structure and provide improved regret over non-causal algorithms. The existing works are based on the hard intervention model, where the causal relations between a node and its parents are completely cut off. Herein, we consider soft interventions, such that the node under intervention can still be causally related to other nodes.
@@
A challenge in MAB problems is the balance between exploration and exploitation. A common approach is via derivations of upper confidence bounds (UCB) [21]. The classic UCB scheme uses the number of visits as a general measure of uncertainty [22], while in causal bandits, uncertainty can be better quantified by bounding the variance of problem-specific estimators [17, 18, 19]. We shall adapt the UCB approach to our context, exploiting properties of our graph identification scheme. Another standard assumption in causal bandits is that of a static causal model. Both [23] and [24] consider the non-stationary case, but assume prior knowledge of the initial causal graph. We shall provide a method by which to keep the causal structure up-to-date without prior graph knowledge.
@@
In this paper, we propose the Causal Sub-graph Learning with Upper Confidence Bound (CSL-UCB) scheme, without assumed knowledge of either the causal graph topology or the interventional distributions. Our algorithm seeks to minimize false negative errors in graph identification as they have a larger impact on optimal intervention design. We quantify the uncertainty in our graph identification coupled with that stemming from decision making. The learning of sub-graphs (versus the entire graph) strongly reduces complexity and provides a natural strategy for change detection in non-stationary environments. The main contributions of this paper are:
@@
    We propose a sub-graph learning approach to learn the causal structure, which has strongly improved computational and sample complexity. A novel mutual information regularization is proposed which reduces false negative errors. It is shown that false negatives have a more critical impact on overall reward optimization in causal bandit problems.
@@
    Numerical comparisons show that the proposed scheme identifies the optimal intervention much faster than standard MAB schemes by exploiting causal structure. Moreover, compared to strategies that only focus on causal graph identification, the proposed scheme has improved sample complexity and learns the causal structure with a limited loss with respect to the long-term reward.
@@ II-A The Causal Graphical Model with Soft Intervention
To represent observational causal effects (no intervention), consider a DAG with structure $(\mathcal{V},\mathcal{B})$, where $\mathcal{V}=\{1,\dots,N\}$ is the set of $N$ nodes and $\mathcal{B}$ is the set of directed edges. Moreover, the observational edge-weight matrix $\bm{B}\in\mathbb{R}^{N\times N}$ captures the strength of causal effects, where the $(i,j)$-th entry represents the weight of the edge $i\rightarrow j$. To model causal effects under intervention, consider node-wise intervention, defined as
$\bm{a}=(a_{1},\dots,a_{N})^{T}\in\{0,1\}^{N},$ (1)
where $a_{i}$ represents whether node $i$ is intervened $(1)$ or not $(0)$. Specifically, instead of hard interventions, we consider soft interventions, which do not cut off causal relationships between the intervened node and its parents, but change the upcoming edges to the node.
Further, we denote the interventional edge-weight matrix by $\bm{B}^{\prime}\in\mathbb{R}^{N\times N}$, such that the post-intervention weight matrix $\bm{B}_{\bm{a}}$ can be constructed as
$\left[\bm{B}_{\bm{a}}\right]_{i}=\mathbb{I}(a_{i}=1)\left[\bm{B}^{\prime}\right]_{i}+\mathbb{I}(a_{i}=0)\left[\bm{B}\right]_{i},$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function and $[\cdot]_{i}$ represents the $i$-th column of a matrix. The $i$-th column of the post-intervention weight matrix determines the set of parents of node $i$ and how these parents causally influence node $i$.
As a result of intervention, the vector of stochastic values associated with the nodes is represented by $\bm{x}\in\mathbb{R}^{N}$. The causal relationship among nodes is described by a linear structure equation model (LinSEM),
$\bm{x}=(\bm{B}_{\bm{a}})^{T}\bm{x}+\bm{\epsilon},$ (3)
where $\bm{\epsilon}$ is a vector of Gaussian noise/exogenous variables, independent of $\bm{x}$. We assume $\bm{\epsilon}$ contains independent elements with known means, denoted by $\bm{\nu}$.
We denote the sets of parents and ancestors of node $i$ by $\mathcal{P}_{i}(a_{i})$ and $\mathcal{A}_{i}(a_{i})$, the estimated set of parents by $\hat{\mathcal{P}}_{i}(a_{i})$. The set difference of the estimated and true parent sets is denoted by $\hat{\mathcal{P}}_{i}\backslash\mathcal{P}_{i}(a_{i})$. Taking node $3$ in Fig. 1 as an example, the parent and ancestor sets are $\mathcal{P}_{3}(0)=\{2\}$, $\mathcal{A}_{3}(0)=\{1,2\}$. If the estimated parent set is $\hat{\mathcal{P}}_{3}(0)=\{2,4\}$, then we have $\hat{\mathcal{P}}_{3}\backslash\mathcal{P}_{3}(0)=\{4\}$.
@@ II-B The Causal Bandit Model
Figure 1: A causal graph with $N=5$ and node $5$ as the reward node.
@@
$\bm{x}=\left(\bm{I}-\bm{B}_{\bm{a}}\right)^{-T}\bm{\epsilon}=\left(\bm{C}_{\bm{a}}\right)^{T}\bm{\epsilon},$ (5)
@@
$\bm{\mu}_{\bm{a}}\doteq\mathbb{E}\left[\left(\bm{I}-\bm{B}_{\bm{a}}\right)^{-T}\bm{\epsilon}\right]=\left(\bm{I}-\bm{B}_{\bm{a}}\right)^{-T}\bm{\nu}.$ (6)
@@ III-A Sub-graph Decomposition
Causal graph identification is equivalent to figuring out causal relationships between every pair of nodes in the graph, represented by directed edges. Given observation up to step $t$, $\bm{X}^{1:t}\doteq(\bm{x}^{1},\dots,\bm{x}^{t})$ of dimension $N\times t$, we can evaluate every possible graph structure by its ability to fit the data. However, finding the best graph structure may not be feasible in practice because the number of possible DAGs grows super-exponentially in the number of nodes. For example, the number of DAGs for $2$, $4$, $10$ nodes are $3$, $543$ and $4.2\times 10^{18}$, respectively [25]. The growing search space issue becomes even more serious for causal bandits, where the agent only has $t$ observations, while it has to identify $2^{N}$ different graphs.
To improve sample efficiency, one feature of the causal bandit model is noteworthy. Even though there are $2^{N}$ post-intervention distributions characterized by $\bm{B}_{\bm{a}}$, those post-intervention weight matrices are composed of columns of $\bm{B}$ and $\bm{B}^{\prime}$. Therefore, instead of identifying $2^{N}$ causal graphs induced by $\bm{B}_{\bm{a}}$, we can identify causal relationships induced by the columns of $\bm{B}$ and $\bm{B}^{\prime}$. Specifically, $\left[\bm{B}\right]_{i}$ and $\left[\bm{B}^{\prime}\right]_{i}$ encode causal weights between node $i$ and its parents. The following proposition ensures that causal graphs can always be uniquely decomposed into sub-graphs of our interest.
@@
Any DAG can always be uniquely decomposed into sub-graphs such that each sub-graph consists of a node and all of its incoming edges.
@@
$\mathcal{V}_{i}=\{i\},\quad\mathcal{B}_{i}=\{(k,i):(k,i)\in\mathcal{B},\forall k\}.$ (9)
Moreover, notice that the decomposition is unique: sub-graph $\mathcal{G}_{i}$ contains and only contains node $i$ and all of its upcoming edges, which completes the proof. ∎
As illustrated in Fig. 1, each sub-graph resides in one of the shaded regions. For example, the sub-graph in the shaded red region consists of node $5$, directed edges $3\rightarrow 5$, $4\rightarrow 5$ and an incoming dashed edge. Learning sub-graphs has exponentially improved data efficiency. Consider an example where interventions are selected in turns up to step $t$, such that each complete graph is associated with $t/2^{N}$ data samples, as there are $2^{N}$ complete graphs. Each sub-graph, instead, is associated with $t/2$ data samples, as each node can only be in one of two possible modes: under intervention, or not.
For learning causal sub-graphs, we employ the principle of independent mechanisms, which states that the causes and the mechanism producing the effect are independent [25]. Based on that, testing the independence of residuals has been investigated for causal inference, in the context of additive noise models [26, 27, 28]. Herein, we consider minimum mean-square error (MMSE) estimation, where the estimator and residual are defined as
$\hat{x}_{i}^{t}(a_{i})\doteq\mathbb{E}\left[x_{i}\middle|\bm{X}^{1:t}_{\hat{\mathcal{P}}_{i}(a_{i})}\right],\quad r_{i}^{t}(a_{i})\doteq\hat{x}_{i}^{t}(a_{i})-x_{i}.$ (10)
With the estimated mechanism producing $x_{i}$, the principle of independent mechanisms implies that the residual should be independent from any causal variable,
@@
Thus, the dependence between $r_{i}^{t}$ and $x_{j}$ supports the hypothesis that the edge $j\rightarrow i$ does not exist. By evaluating all potential in-coming edges to a node, the corresponding sub-graph can be estimated.

### v2

@@ Abstract
In this paper, the causal bandit problem is investigated, with the objective of maximizing the long-term reward by selecting an optimal sequence of interventions on nodes in an unknown causal graph. It is assumed that both the causal topology and the distribution of interventions are unknown. First, based on the difference between the two types of graph identification errors (false positives and negatives), a causal graph learning method is proposed. Numerical results suggest that this method has a much lower sample complexity relative to the prior art by learning sub-graphs. However, we note that a sample complexity analysis for the new algorithm has not been undertaken, as of yet. Under the assumption of minimum-mean squared error weight estimation, a new uncertainty bound tailored to the causal bandit problem is derived. This uncertainty bound drives an upper confidence bound-based intervention selection to optimize the reward. Further, we consider a particular instance of non-stationary bandits wherein both the causal topology and interventional distributions can change. Our solution is the design of a sub-graph change detection mechanism that requires a modest number of samples. Numerical results compare the new methodology to existing schemes and show a substantial performance improvement in stationary and non-stationary settings. Averaged over $100$ randomly generated causal bandits, the proposed scheme takes significantly fewer samples to learn the causal structure and achieves a reward gain of $85\%$ compared to existing approaches.
@@
Causal bandit, linear structural equation model, graph identification, upper confidence bound.
@@ I Introduction
Investigation of cause-effect relationships is central to answering key questions in medicine, epidemiology, economics as well as the social and behavioral sciences (see e.g. [2, 3, 4]). Cause-effect relationships can be represented by Bayesian networks, in the form of directed acyclic graphs (DAGs). Identifying the causal structural model is essential to decision making and answering counterfactual questions. Herein, we will study methods for determining the underlying causal structural model in the context of causal bandits and thus will also optimize intervention design. We first discuss causal graph identification and then discuss causal bandits.
I-A Related Work
@@
Current causal graph identification methods mainly fall into two categories: constraint-based and score-based. Constraint-based approaches, such as[2, 5], search for conditional independence over all possible graphs and thus computation complexity scales exponentially. Classical score-based methods evaluate different graph structures in their ability to fit the data (see e.g. [6, 7, 8]). To lower the complexity, recent score-based methods (see e.g. [9, 10, 11, 12]) formulate the structure learning problem as a continuous optimization task. In [10], an algorithm called GOLEM is proposed, where the likelihood score is combined with a soft DAG constraint to evaluate possible causal structures. The DAGMA scheme proposed in [11] characterizes acyclicity by a log-determinant function, with better-behaved gradients.
@@
Decision-making or action-taking is natural within the context of causal inference. Despite this natural relationship, incorporation of causal methods within the context of multi-armed bandits (MAB) [13, 14], has only occurred recently. MABs provide a useful model for sequential decision-making problems, and as such have been employed in the design of clinical trials [15], recommendation systems [16], financial portfolio design [17], etc. In a MAB, an agent selects an action (arm) in each round and observes corresponding outcomes – the goal is to maximize the long-term cumulative reward. In the standard setting, the stochastic rewards generated by different arms are assumed to be statistically independent.
To model realistic scenarios with dependence, structured bandits have been considered. Within this arena, we focus on causal bandits [14], where the causal structure is exploited to improve decision-making. DAGs can encode the causal relationship among factors that contribute to the reward [14]. Herein, we interpret the arms as different interventions on the nodes of a DAG and the reward as the stochastic outcome of a certain node. In the current work, we will provide new methods for both causal discovery (graph identification) as well as the design of optimal interventions for causal MABs.
TABLE I: Key Assumptions: Related Work
Reference Topology Knowledge Intervention Structural Model Non-stationary Structure Exogenous/Noise
[14, 18, 19] Hard Discrete Bernoulli
[20] Soft Linear Sub-Gaussian
[21] Soft General Known distribution
[22] Soft General Sub-Gaussian
[23] Known Separating Sets Hard Discrete & Linear Bernoulli & Gaussian
[24] Hard Discrete Sub-Gaussian
[25] Hard Linear sub-Gaussian
[26] Soft Linear sub-Gaussian
This work Soft Linear Gaussian
@@
Existing literature on causal bandits can be categorized based on their assumptions about the topology of the underlying DAG and the distribution of the interventions. While many works assume prior topology or DAG knowledge [14, 18, 19, 21, 20, 22], this is often not known in practice. In this work, we assume that the underlying causal structure is unknown.
@@
The setting without the knowledge of both topology and interventional distribution has been investigated recently [24, 23, 25, 27]. An algorithm based on causal tree recovery is proposed in [24], with regret scaling logarithmically with the number of nodes. In [23], an auxiliary separating set algorithm is utilized to learn the causal structure and provide improved regret over non-causal algorithms. The existing works are based on the hard (perfect) intervention model, where the causal relations between a node and its parents are completely cut off because a specific value is assigned to the node. Herein, we consider soft interventions, such that the node under intervention can still be causally related to other nodes.
@@
A challenge in MAB problems is the balance between exploration and exploitation. A common approach is via derivations of upper confidence bounds (UCB) [28]. The classic UCB scheme uses the number of visits as a general measure of uncertainty [29], while in causal bandits, uncertainty can be better quantified by bounding the variance of problem-specific estimators [20, 24, 23]. We shall adapt the UCB approach to our context, exploiting properties of our graph identification scheme. Another standard assumption in causal bandits is that of a static causal model. Both [26] and [30] consider the non-stationary case, but assume prior knowledge of the initial causal graph. We shall provide a method by which to keep the causal structure up-to-date without prior graph knowledge. A comparison between prior works and our approach is provided in Table I.
I-B Approach and Contributions
@@
We note that separating graph identification from optimized interventions for MABs is inefficient as these two goals are coupled in causal MABs. Another feature not typically considered in the causal graph identification or causal MAB literature is the distinguishing between false positive and false negative errors in graph identification – we shall actively consider this feature. A false positive means that the graph identification method outputs an edge that is not present in the true graph, while a false negative means that the graph identification method outputs no edge, where in fact, the edge is present in the true DAG. In this paper, we propose the Causal Sub-graph Learning with Upper Confidence Bound (CSL-UCB) scheme, without assumed knowledge of either the causal graph topology or the interventional distributions. Our algorithm seeks to minimize false negative errors in graph identification as they have a larger impact on optimal intervention design. We quantify the uncertainty in our graph identification coupled with that stemming from decision making. The learning of sub-graphs (versus the entire graph) strongly reduces complexity and provides a natural strategy for change detection in non-stationary environments. The main contributions of this paper are:
@@
    We propose a sub-graph learning approach to learn the causal structure, which has strongly improved computational complexity. Numerical results suggest that it has a much lower sample complexity as well. A novel mutual information regularization is proposed which reduces false negative errors. It is shown that false negatives have a more critical impact on overall reward optimization in causal bandit problems.
@@
    Numerical comparisons show that the proposed scheme identifies the optimal intervention much faster than standard MAB schemes by exploiting causal structure. Moreover, compared to strategies that only focus on causal graph identification, the proposed scheme needs much fewer samples than other methods and learns the causal structure with a limited loss with respect to the long-term reward.
@@ II-A The Causal Graphical Model with Soft Intervention
We assume the causal effects can be captured by a DAG with structure $(\mathcal{V},\mathcal{B})$, where $\mathcal{V}=\{1,\dots,N\}$ is the set of $N$ nodes and $\mathcal{B}$ is the set of directed edges. Moreover, the edge-weight matrix $\bm{B}\in\mathbb{R}^{N\times N}$ captures the strength of causal effects, where the $(i,j)$-th entry represents the weight of the edge $i\rightarrow j$. To describe observations under intervention, consider node-wise intervention, defined as
$\bm{a}={(a_{1},\dots,a_{N})}^{\top}\in\{0,1\}^{N},$ (1)
where $a_{i}$ represents whether node $i$ is intervened ($a_{i}=1$) or not ($a_{i}=0$). Specifically, instead of hard interventions, we consider soft interventions, which do not necessarily cut off causal relationships between the intervened node and its parents, but change the in-coming edges to the node. To formalize the model for soft intervention, we denote the interventional edge-weight matrix by $\bm{B}^{\prime}\in\mathbb{R}^{N\times N}$, such that the post-intervention weight matrix $\bm{B}_{\bm{a}}$ can be constructed as
$\left[\bm{B}_{\bm{a}}\right]_{i}=\mathbb{I}(a_{i}=1)\bm{B}^{\prime}_{i}+\mathbb{I}(a_{i}=0)\bm{B}_{i},$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function and $[\cdot]_{i}$ represents the $i$-th column of a matrix such that $\bm{B}_{i},\bm{B}^{\prime}_{i}\in\mathbb{R}^{N\times 1}$. The $i$-th column of the post-interventional weight matrix determines the set of parents of node $i$ and how these parents causally influence node $i$.
As a result of intervention, the vector of stochastic values associated with the nodes is represented by $\bm{x}\in\mathbb{R}^{N}$. The causal relationship among nodes is described by a linear structural equation model (LinSEM),
$\bm{x}={(\bm{B}_{\bm{a}})}^{\top}\bm{x}+\bm{\epsilon},$ (3)
where $\bm{\epsilon}$ is a vector of Gaussian exogenous/noise variables. We assume $\bm{\epsilon}$ contains independent elements with known means $\bm{\nu}$ and unknown variances $\bm{\sigma}^{2}$. Further, we denote $\tilde{\bm{\epsilon}}\doteq\bm{\epsilon}-\bm{\nu}$ such that $\tilde{\bm{\epsilon}}\sim\mathcal{N}(\bm{0},\mathrm{diag}(\bm{\sigma}^{2}))$.
We denote the sets of parents and ancestors of node $i$ by $\mathcal{P}_{i}(a_{i})$ and $\mathcal{A}_{i}(a_{i})$, the estimated set of parents by $\widehat{\mathcal{P}}_{i}(a_{i})$. The set difference of the estimated and true parent sets is denoted by $\widehat{\mathcal{P}}_{i}\backslash\mathcal{P}_{i}(a_{i})$. Taking node $3$ in Fig. 1 as an example, the parent and ancestor sets are $\mathcal{P}_{3}(0)=\{2\}$, $\mathcal{A}_{3}(0)=\{1,2\}$. If the estimated parent set is $\widehat{\mathcal{P}}_{3}(0)=\{2,4\}$, then we have $\widehat{\mathcal{P}}_{3}\backslash\mathcal{P}_{3}(0)=\{4\}$.
@@ II-B The Causal Bandit Model
Figure 1: A causal graph with $N=5$ and node $5$ as the reward node. Each sub-graph is marked by a colored region. The sets of parents and ancestors are provided in the inline table.
@@
$\bm{x}={\left(\bm{I}-\bm{B}_{\bm{a}}\right)}^{-\top}\bm{\epsilon}={\left(\bm{C}_{\bm{a}}\right)}^{\top}\bm{\epsilon},$ (5)
@@
$\bm{\mu}_{\bm{a}}\doteq\mathbb{E}\left[{\left(\bm{I}-\bm{B}_{\bm{a}}\right)}^{-\top}\bm{\epsilon}\right]={\left(\bm{I}-\bm{B}_{\bm{a}}\right)}^{-\top}\bm{\nu}.$ (6)
@@ III-A Sub-graph Decomposition
Causal graph identification is equivalent to figuring out causal relationships between every pair of nodes in the graph, represented by directed edges. Given observation up to step $t$, $\bm{X}^{1:t}\doteq(\bm{x}^{1},\dots,\bm{x}^{t})$ of dimension $N\times t$, we can evaluate every possible graph structure by its ability to fit the data. However, finding the best graph structure may not be feasible in practice because the number of possible DAGs grows super-exponentially in the number of nodes. For example, the number of DAGs for $2$, $4$, $10$ nodes are $3$, $543$ and $4.2\times 10^{18}$, respectively [31]. The growing search space issue becomes even more serious for causal bandits. Since the effect of an intervention is characterized by the post-interventional graph $\bm{B}_{\bm{a}}$ and there are $2^{N}$ possible interventions, the agent has to identify $2^{N}$ different graphs with only $t$ observations.
To improve sample efficiency, one feature of the causal bandit model is noteworthy. Even though there are $2^{N}$ post-intervention distributions characterized by $\bm{B}_{\bm{a}}$, those post-intervention weight matrices are composed of columns of $\bm{B}$ and $\bm{B}^{\prime}$ [20]. Therefore, instead of identifying $2^{N}$ causal graphs induced by $\bm{B}_{\bm{a}}$, we can identify causal relationships induced by the columns of $\bm{B}$ and $\bm{B}^{\prime}$. Specifically, $\bm{B}_{i}$ and $\bm{B}^{\prime}_{i}$ encode causal weights between node $i$ and its parents. The following proposition ensures that causal graphs can always be uniquely decomposed into sub-graphs of our interest. [^1] The sub-graphs we consider here are not vertex-induced sub-graphs.
@@
Any DAG can always be uniquely decomposed into sub-graphs such that each sub-graph consists of a node and all in-coming edges to the node.
@@
$\mathcal{V}_{i}=\{i\},\quad\mathcal{B}_{i}=\{(k,i):\forall k\in\mathcal{P}_{i}\}.$ (9)
Moreover, notice that the decomposition is unique: sub-graph $\mathcal{G}_{i}$ contains and only contains node $i$ and all of its in-coming edges, which completes the proof. ∎
As illustrated in Fig. 1, each sub-graph resides in one of the colored regions. For example, the sub-graph in the red region consists of node $5$, directed edges $(3,5)$ and $(4,5)$. Learning sub-graphs has exponentially improved data efficiency. Consider an example where interventions are selected in turns up to step $t$, such that each complete graph is associated with $t/2^{N}$ data samples, as there are $2^{N}$ complete graphs. Each sub-graph, instead, is associated with $t/2$ data samples, as each node can only be in one of two possible modes: under intervention, or not.
For learning causal sub-graphs, we employ the principle of independent mechanisms, which states that the causes and the mechanism producing the effect are independent [31]. Based on that, testing the independence of residuals has been investigated for causal discovery, in the context of additive noise models [32, 33, 34]. Herein, we consider minimum mean-square error (MMSE) estimation, where the estimator and residual are defined as
$\hat{x}_{i}^{t}(a_{i})\doteq\mathbb{E}\left[x_{i}\middle|\bm{X}^{1:t}_{\widehat{\mathcal{P}}_{i}(a_{i})}\right],\quad r_{i}^{t}(a_{i})\doteq\hat{x}_{i}^{t}(a_{i})-x_{i}.$ (10)
With the estimated mechanism producing $x_{i}$, the principle of independent mechanisms implies that the residual should be independent from values of the parent nodes,
@@
Thus, the dependence between $r_{i}^{t}$ and $x_{j}$ supports the hypothesis that node $j$ is not a parent of node $i$ and thus the edge $j\rightarrow i$ does not exist. By evaluating all potential in-coming edges to a node, the corresponding sub-graph can be estimated.